A square hall is surrounded by a veranda of width 2.5m on its outside. If area of veranda is 175m2 , what is the area of the hall
Answer by ankor@dixie-net.com(22740)

A square hall is surrounded by a veranda of width 2.5m on its outside.
If area of veranda is 175m2 , what is the area of the hall
:
let x = the length of the square hall
then
x^2 = the area of the hall
:
The 2.5 m veranda around the outside adds 5 m to both overall dimensions
(x+5) by (x+5) = the overall dimensions
:
Overall area - hall area = 175 sq/m
(x+5)(x+5) - x^2 = 175
FOIL
x^2 + 10x + 25 - x^2 = 175
combine like terms
10x = 175 - 25
10x = 150
x = 150/10
x = 15 m length of the hall
there
15^2 = 225 sq/m area of the hall
:
:
:
Check this by finding the areas
20^2 = 400
15^2 = 225
------------
diff: 175